OpenStudy (anonymous):

Add: (3x4 + x2 + 2) + (9x4 - x2 + x)

OpenStudy (anonymous):

think of this as "combine like terms"

OpenStudy (anonymous):

but is it x4

pooja195 (pooja195):

\[\huge~\rm~ 3x^4 + x^2 + 2 + 9x^4 - x^2 + x\]

OpenStudy (anonymous):

\[3x^4+9x^4=?\]

OpenStudy (anonymous):

so it's 12x4?

OpenStudy (anonymous):

the first term is, yes

OpenStudy (anonymous):

then \[x^2-x^2=?\]

OpenStudy (anonymous):

ok the second term is x2

OpenStudy (anonymous):

hmm no

OpenStudy (anonymous):

what is \[x^2-x^2\]?

OpenStudy (anonymous):

ok

pooja195 (pooja195):

Think of it as \[\huge~\rm~1x^2−1x^2\] 1-1=?

OpenStudy (anonymous):

that's 0 so the answer is x2 but you guys said no so i'm not sure

pooja195 (pooja195):

They cancel out

OpenStudy (anonymous):

so war does that equal

pooja195 (pooja195):

0

pooja195 (pooja195):

Therefore we ignore it

OpenStudy (anonymous):

and then it's x2 right?

pooja195 (pooja195):

\[\huge~\rm~(3x^4+9x^4)+(x^2+−x^2)+(x)+(2)\] Ok so we have 12x^4+x+2

OpenStudy (anonymous):

none of those are like terms so is that the answer

OpenStudy (anonymous):

right?

pooja195 (pooja195):

yep

pooja195 (pooja195):

12x^4+x+2

OpenStudy (anonymous):

ok thanks a lot i don't know who to give the medal to